OpenStudy (darkigloo):

Calculus: Force..A dam is inclined at an angle of 30∘ from the vertical and has the shape of an isosceles trapezoid 100 ft wide at the top and 50 ft wide at the bottom and with a slant height of 70 ft. Find the hydrostatic force (in lb) on the dam when it is full of water.

OpenStudy (xapproachesinfinity):

wanna how is this done

OpenStudy (darkigloo):

all i know is that the perpendicular height h is 70cos(30)= 60.6 ft

OpenStudy (darkigloo):

yes

OpenStudy (xapproachesinfinity):

so the force on the dam is directed where?

OpenStudy (darkigloo):

the triangle?

OpenStudy (xapproachesinfinity):

what triangle do you mean?

OpenStudy (darkigloo):

oh im sorry, i was thinking of another question. i don't know where its being directed.

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

actually this seems much good to me

OpenStudy (darkigloo):

thats how i drew it.

OpenStudy (xapproachesinfinity):

ok so what's the problem

OpenStudy (darkigloo):

not sure how to approach the problem

OpenStudy (xapproachesinfinity):

what do you know about hydrostatic forces

OpenStudy (xapproachesinfinity):

that could be the link your looking for the fact that you have all the given looks like this is related area

OpenStudy (xapproachesinfinity):

link to what you are looking for*

OpenStudy (dumbcow):

First you need formula for finding pressure and Force \[P = \rho g d\] \[\rho = 62.4 lb/ft^3, g = 32 ft/s^2\] d = depth under water \[F = PA\] where A is surface Area  The total force of trapezoid will be the sum of all the rectangle forces of height "dy" \[A = 2x dy = 2(\frac{25}{70}y +25) dy = (\frac{5}{7} y + 50) dy\] depth is the perpendicular height \[d = \cos 30 * y = \frac{\sqrt{3}}{2} y\] Finally put it all together to get Force equation, then integrate over "y" from 0 to 70 \[F = \int\limits_0^{70}(62.4)(32)(\frac{\sqrt{3}}{2}y)(\frac{5}{7}y+50) dy\] \[F = (62.4)(16 \sqrt{3}) \int\limits_0^{70} [\frac{5}{7} y^2 + 50y] dy\] Integrate using power rule, apply limits and you are done